OpenStudy (anonymous):

Write 4420 as a sum of 2 squares? (No programs, use 2 square identity)

OpenStudy (anonymous):

I can't solve this for you, as my maths isn't good enough, but I can suggest a start: A number ending in 0 squared is a number ending in 00. A number ending in 1 or 9 squared is a number ending in 1. A number ending in 2 or 8 squared is a number ending in 4. A number ending in 3 or 7 squared is a number ending in 9. A number ending in 4 or 6 squared is a number ending in 6. A number ending in 5 squared is a number ending in 25. Thus you can eliminate some possibilities for the squares you search for.

OpenStudy (anonymous):

@Cheese3.1416 Well, that's OK, a computer program might use this to narrow things down. But it is simpler to use the two square identity (a^2+b^2)(c^2+d^2) = (ac+bd)^2 + (ad-bc)^2

OpenStudy (anonymous):

aka the Brahmagupta–Fibonacci identity

OpenStudy (anonymous):

You seem to know what you're doing better than I do. Is this a test?

OpenStudy (anonymous):

Think of it as a tutorial...

OpenStudy (anonymous):

OK, challenge accepted. Whether it'll be completed is entirely another matter....

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Maybe this is me being stupid, but all I can say about the two square identity is it gives me one equation with 4 unknowns.

OpenStudy (shubhamsrg):

*

OpenStudy (anonymous):

OK, a hint then: 5 is a factor of 4420 and you could write that as (2^2 +1^2)

OpenStudy (anonymous):

So now the problem becomes write 884 as the sum of two squares?

OpenStudy (anonymous):

Well, yes, 5 is not the only factor....

OpenStudy (anonymous):

This is starting to look recursive...

OpenStudy (anonymous):

First off, you better get a factorization of 4420

OpenStudy (anonymous):

\[2 \times 2 \times 5 \times 221\]

OpenStudy (anonymous):

\[221=14^2+5^2\]

OpenStudy (anonymous):

Yes, u could have bro0ken 221 into 13 *17

OpenStudy (anonymous):

\[20=2^2+4^2\]

OpenStudy (anonymous):

There is more than 1 answer so let's use 2^2 * 5 * 13 * 17 as the factorization

OpenStudy (anonymous):

I really need to make myself more intelligent.

OpenStudy (anonymous):

That is just the usual factorization...

OpenStudy (anonymous):

Small numbers , easy to put as sums of squares...

OpenStudy (anonymous):

I am now processing.

OpenStudy (anonymous):

Save a little time (2^2+0^2) (2^2+1^2) (3^2+2^2) (4^2+1^2)

OpenStudy (anonymous):

There was no need to give me the answer.

OpenStudy (anonymous):

You spoiler.

OpenStudy (anonymous):

That's not the answer....

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Well I can work it out myself.

OpenStudy (anonymous):

I'm about to check 130 and 34.

OpenStudy (anonymous):

But u don't have to check anything, just use the 2 square identity....

OpenStudy (anonymous):

Ok, well if\[130=11^2+3^2\]\[34=5^2+3^2\]then I guess\[(33+15)^2+(33-15)^2=4420\]

OpenStudy (anonymous):

Try again...

OpenStudy (anonymous):

Eh, stupid me. I should've realized instantly that's wrong.

OpenStudy (anonymous):

Right idea, sort of...

OpenStudy (anonymous):

More systematic wouold be to apply the identity 3 times to (2^2+0^2) (2^2+1^2) (3^2+2^2) (4^2+1^2)

OpenStudy (anonymous):

I'm no good at being systematic.

OpenStudy (anonymous):

I will give you a medal for effort (and negative brownie points for being unsystematic)

OpenStudy (anonymous):

That seems generous.

OpenStudy (anonymous):

Damn. The moment I said this problem seemed recursive I should've realized I needed to do what you said 4 posts ago.

OpenStudy (anonymous):

Yes, procedure is simply a recursive algorithm when all is said and done...

OpenStudy (anonymous):

Thanks for the medal, though I personally feel a bit disappointed with myself.

OpenStudy (anonymous):

S'OK, it's good practice...